DIRECT OBSERVATION OF THE 2 ^{3}Π_{u} STATE OF Rb_{2} IN A PULSED MOLECULAR BEAM: ROTATIONAL-BRANCH INTENSITY ANOMALIES IN THE 2 ^{3}Π_{u}(1_{u}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) BANDS

\underline{Y. LEE},Y. YOON, S. J. BAEK, D-L JOO, J-S RYU and B. KIM, Department of Chemistry, Korea Advanced Institute of Science and Technology, Taejon 305-701, Korea.

The first observation of 2 ^{3}Π_{u} − X^{1}Σ_{g}^{+} transitions is reported. Rotationally resolved transitions of the 2 ^{3}Π_{u}(1_{u}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) and 2 ^{3}Π_{u}(0_{u}^{+}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) are observed by resonance enhanced 2-photon ionization (RE2PI) method in a pulsed molecular beam. Ω-doubling and interference induced rotational branch intensity anomalies are observed for 2 3Π_{u}(1_{u}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) transitions. Both of 2 ^{3}Π_{u}(0_{u}^{+}) and 2 ^{3}Π_{u}(1_{u}) states are strongly mixed with singlet states by spin-orbit coupling. The former with 2 ^{1}Σ_{u}^{+} and the latter with 2 ^{1}Π_{u}. In relatively weak 2 ^{3}Π_{u}(1_{u}) bands P −branch rotational lines disappear and the intensities of R−branch rotational lines are enhanced. These intensity anomalies in 2 3Π_{u}(1_{u}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) transition, due to an interference effect between parallel and perpendicular transition amplitudes, is caused by ΔΩ = ±1 perturbation. The molecular constants of 2 ^{3}Π_{u}(0_{u}^{+}) − X^{1}Σ_{g}^{+}(0_{g}^{+}) transition are determined as T_{e} = 19784.2588 ± 0.0088 cm^{−1}, ω_{e} = 42.1954 ± 0.0060 cm^{−1}, ω_{e}x_{e} = 0.1701 ± 0.0011 cm^{−1}, ω_{e}y_{e} = −0.001096 ± 0.000057 cm^{−1}, B_{e} = 0.018503±0.000018 cm^{−1} for ^{8}5Rb_{2}. The spin−orbit coupling constant A and Ω-doubling parameters p and q are determined by simultaneous fitting of the rotational contours of both 2^{3}Π_{u}(1_{u})−X^{1}Σ_{g}^{+}(0_{g}^{+}) and 2^{3}Π_{u}(0_{u}^{+})−X^{1}Σ_{g}^{+}(0_{g}^{+}) transitions.